OpenStudy (anonymous):

Find the common difference of the sequence shown. 1/6, 1/3, 1/2, 2/3... A. 1/2 B. 1/6 C. 1/3

OpenStudy (anonymous):

give them all the same denominator: \[\frac{ 1}{ 6 }, \frac{ 2}{ 6 },\frac{ 3}{ 6 },\frac{ 4}{ 6 }\] Make a little more sense?

OpenStudy (anonymous):

YEA 1/6 Thanks!!